1, If there is a 20% chance we will get a -2% return, a 30% chance of getting a 6% return, and a 50% chance of getting 17% return. What is the expected rate of return for the investment?

Please round your answer to the fourth decimal. If your answer is 1.11%, then you should input "0.0111".

2, Assume your grandparents deposit $268,402 into a bank account for you and will allow you to withdraw $71,275 at the end of every year, starting the end of this year. The bank pays an annual interest rate of 4.75% compounded annually. How many years can you withdraw the money and leave the account completely empty after the last withdraw?

Please round your answer to the second decimal. E.g. 1.50 (years).
Answer #1

1.Expected return=Respective return*Respective probability

=(0.2*-2)+(0.3*6)+(0.5*17)

=0.099

2.Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

268,402=71,275[1-(1.0475)^-time period]/0.0475

268,402=1500526.32[1-(1.0475)^-time period]

(268,402/1500526.32)=1-(1.0475)^-time period

(1.0475)^-time period=1-(268,402/1500526.32)

(1/1.0475)^time period=0.821128096

Taking log on both sides;

time period*log(1/1.0475)=log0.821128096

time period=log 0.821128096/log(1/1.0475)

=4.25 years(Approx)